A certain metal is 20% tin. How many kilograms of this metal must be mixed with 80 kilograms of a metal that is 70% tin to get a metal that is 50% tin?
Answer by sachi(548)

let x kg of 20% mixed with 80 kg of 70% to get 50%
so x*20%+80*70%=(x+80)*50%
or 20x+5600=50x+4000
or 30x=1600
or x=1600/30=53.33 kg
ans